OpenStudy (anonymous):

Multiply and simplify: b4 • b • b2 • b3 You could save my grade, i'd appreciate help :(

OpenStudy (the_fizicx99):

b4, b2 and b3... Are they \[b^4 \times b \times b^2 \times b^3\] ?

OpenStudy (the_fizicx99):

Anyways, b^4 in expanded form is b * b * b * b When you multiply a term with the same base you add the exponents, \(\ a^2 * a = a^{2+1} = a^3 \) \(\ a = a^1 \) So \(\ b^4 * b * b^2 * b^3\) is equal to \(\ b^{4 + 1 + 2 + 3}\) = \(\ b^{10}\)

OpenStudy (anonymous):

Wow, thank you so much! You explained it clearly and I understood. Thanks!!!

OpenStudy (the_fizicx99):

Yw :>